OpenStudy (anonymous):

A portion of a parabola is defined by the parametric equations x = 2 − (t + 1)^2, y = t + 2 with −3 ≤ t ≤ 0. (a) Eliminate the parameter t to determine the equation of the curve in terms of only x and y. (b)sketch the graph of the parametric curve only for −3 ≤ t ≤ 0. Use arrows to indicate the direction of increasing t.

OpenStudy (anonymous):

Clearly:\[ y=t+2\implies y-2= t \]Thus we get: \[ x = 2 − (t + 1)^2 \implies x = 2 − ((y-2)+ 1)^2 \]